OpenStudy (anonymous):

PLEASE HELP!!!!! A right triangle has base (x - 7) units and height (2x - 10) units. Part A: What is the square of the length of the hypotenuse of the triangle? Show your work. (4 points) Part B: What is the area of the triangle? (3 points) Part C: Using the solution obtained in Part B, explain the closure property of multiplication of polynomials. (3 points)

OpenStudy (anonymous):

Part A i believe is this: hypotenuse^2 = (x-7)^2 +(2x -10)^2

OpenStudy (welshfella):

part A this is the pythagoras theorem

OpenStudy (anonymous):

Right, i know =) ^^^^^

OpenStudy (anonymous):

I need help with Part B and C really =)

OpenStudy (welshfella):

the area of the triangle = 0.5*base*height

OpenStudy (anonymous):

Okay so. . . . . . I need a little bit of help with that pls =)

OpenStudy (anonymous):

is it x^2 - 12x +35?

OpenStudy (welshfella):

0.5(x - 7(2x - 10) distribute the x and -7 over the second parentheses first then do the multiplication by 0.5

OpenStudy (anonymous):

@welshfella

OpenStudy (welshfella):

0.5(x^2 - 24x + 70)

OpenStudy (welshfella):

your almost right

OpenStudy (welshfella):

as for part C - the result is a polynomial which demonstrates that polynomials are closed under multiplication. Any polynomial times another gives a polynomial.

OpenStudy (anonymous):

Okay so part B is 0.5x^2 -12x +35 ?

OpenStudy (welshfella):

yes

OpenStudy (anonymous):

okay now for part C =)

OpenStudy (welshfella):

oops sorry !! you didn't spot my deliberate mistake you were right first time its 0.5(2x^2 - 24x + 10) = x^2 - 12x + 35

OpenStudy (welshfella):

I've answered part C already.

OpenStudy (anonymous):

Oh okay so it's x^2 -12x +35?

OpenStudy (anonymous):

Part C: Well the closure property of polynomials is that if you multiply a polynomial by a polynomial you will end up with a polynomial ?

OpenStudy (welshfella):

yes

OpenStudy (anonymous):

Part C: Well the closure property of polynomials is that if you multiply a polynomial by a polynomial you will end up with a polynomial. So if we multiply x^2 -12x +35 by it’s self we will get a polynomial.

OpenStudy (welshfella):

we sure will

OpenStudy (anonymous):

So do u think that will answer part C sufficiently? =)

OpenStudy (welshfella):

yes

OpenStudy (welshfella):

its quite simple really any polynomial multiplied by itself or another polynomial will always give another polynomial. So its closed under multiplication.

OpenStudy (welshfella):

just as integers are closed under multiplication . The result will always be an integer.

OpenStudy (welshfella):

i'm not sure if they want you to expand and simplify the expression for the square of the hypotenuse. What do you think? Its worth 4 points.

OpenStudy (anonymous):

Okay awesome thx! =)

OpenStudy (welshfella):

@CountryGurl15